# Building 3D In-Context Learning Universal Model in Neuroimaging

### v1

@@ Abstract
In-context learning (ICL), a type of universal model, demonstrates exceptional generalization across a wide range of tasks without retraining by leveraging task-specific guidance from context, making it particularly effective for the complex demands of neuroimaging. However, existing ICL models, which take 2D images as input, struggle to fully leverage the 3D anatomical structures in neuroimages, leading to a lack of global awareness and suboptimal performance. In this regard, we introduce Neuroverse3D, an ICL model capable of performing multiple neuroimaging tasks (e.g., segmentation, denoising, inpainting) in 3D. Neuroverse3D overcomes the large memory consumption due to 3D inputs through adaptive parallel-sequential context processing and a U-shape fusion strategy, allowing it to handle an unlimited number of context images. Additionally, we propose an optimized loss to balance multi-task training and enhance the focus on anatomical structures. Our study incorporates 43,674 3D scans from 19 neuroimaging datasets and evaluates Neuroverse3D on 14 diverse tasks using held-out test sets. The results demonstrate that Neuroverse3D significantly outperforms existing ICL models and closely matches the performance of task-specific models. The code and model weights are publicly released at: https://github.com/jiesihu/Neu3D.
@@ 1 Introduction
Computational neuroimage analysis has significantly advanced brain understanding and non-invasive diagnostics, crucial for quantitative analysis and precision medicine. Recently, universal models trained on multi-domain datasets have gained attention for their ability to handle diverse tasks (e.g., segmentation, denoising, inpainting) and modalities (e.g., T1, T2, CT), demonstrating adaptability to domain shifts and even unseen tasks [11, 36, 41, 55]. In-context learning (ICL) has emerged as a particularly promising paradigm [brown2020language, 49, 7]. ICL models rapidly adapt to specific domains without retraining by utilizing image-label pairs as prompts to provide task-specific guidance. These informative prompts make ICL models particularly effective for segmenting tissues with complex morphology and unifying diverse image generation tasks.
@@
While current ICL models demonstrating clinical potential have been applied to neuroimaging [11, 13, 52], they suffer from inherent dimensional limitations. Existing ICL models process 3D volumes through 2D slices, discarding crucial inter-slice correlations and volumetric context. This lack of global spatial awareness makes them suboptimal for neuroimaging tasks requiring complex anatomical understanding.
@@
To address these limitations, developing an ICL model that captures 3D global information is essential. However, creating a universal ICL model with 3D input presents significant challenges, as 3D images can be over 100 times larger than their 2D slices, leading to substantial memory requirements. This results in three critical issues: (1) Context size, a key factor for ICL performance, becomes severely limited. Larger context sizes have been shown to significantly improve performance [11, 7, 13, 50]. (2) Training models with more parameters becomes infeasible, restricting their capacity and potential performance. (3) Processing high-resolution 3D images becomes difficult.
@@
Figure 1: (a) Parallel processing of all contexts, which is applied by most existing ICL models. (b) Sequential processing of each context, which greatly reduces memory usage. (c) Shows the memory increment when increasing the context size. Our proposed model processes contexts in an adaptive mode, allowing for flexible parallel and sequential processing of contexts.
@@
To develop 3D universal models in neuroimaging, we introduce Neuroverse3D, an in-context learning model take 3D neuroimaging as input. To our knowledge, it is the first 3D ICL universal model for medical imaging. To mitigate the significant memory demands, we propose Adaptive Parallel-Sequential Processing (APSP) to partition the entire context into multiple mini-contexts for processing, while ensuring the consistency of the model’s gradient when training and output when inferencing. The proposed U-shape fusion strategy ensures the efficient handling of APSP. As shown in Figure 1, this approach can significantly reduce memory consumption, even accommodating unlimited context sizes. Additionally, to address the more severe task imbalance issues arising from increased dimensionality and enhance the focus on brain anatomy, we propose an optimized loss function. As the first 3D ICL model in neuroimaging, we evaluate Neuroverse3D’s potential across diverse tasks on held-out datasets. Our contributions are summarized as follows:
  • We propose Neuroverse3D, the first ICL universal model for 3D neuroimaging, trained on extensive multi-center datasets, which provides valuable insights for the development of universal models in neuroimaging.
  • We introduce a novel APSP approach with a U-shaped fusion strategy, allowing ICL models to adaptively balance parallel and sequential context processing based on computational resources. This effectively overcomes memory constraints associated with large context inputs.
@@
  • We develop an optimized loss function to address class imbalance resulting from increased dimensionality and enhance focus on challenging segmentation regions and brain anatomy, further improving performance.
  • We evaluate Neuroverse3D on held-out datasets, demonstrating substantial improvements over other ICL models, with gains exceeding 20 absolute points in Dice on four tasks, and performance on par with task-specific models.
@@ 2.1 Domain Challenges in Medical Imaging
Deep learning models in medical imaging frequently face domain shifts due to heterogeneous imaging data distributions, which reduce model performance on new, unseen domains. Solutions like domain adaptation [57, 53, 27, 51] require target-domain fine-tuning, limiting practical use due to the need for deep learning expertise. Domain generalization aims to produce models that generalize to new domains without fine-tuning [28, 30, 56, 47]. While it alleviates the effects of domain shift, generalization remains challenging as models lack knowledge about unseen domains during training. Different from domain methods, universal models have emerged as a promising solution by learning generalizable image representations from a large number of datasets.
@@ 2.2 Universal Models in Medical Imaging
Universal models in medical imaging can be classified into three categories based on their prompt types. The first category utilizes symbolic prompts, such as points or bounding boxes, as demonstrated by models like SAM [41, 10, 32]. The second category employs natural language prompts [36, 37]. The third category, also known as ICL models, uses image-label pairs as prompts [13, 11, 52, 26]. These universal models, trained on extensive datasets, exhibit exceptional generalization capabilities. Among these, image-label pairs provide rich contextual information, enabling ICL models to perform accurate segmentation on unseen tasks and even handle image generation tasks. However, the richness of these prompts in ICL models’ inputs significantly increases memory consumption, a critical challenge this study seeks to overcome.
@@ 2.3 In-Context Learning
Originally developed in natural language processing [brown2020language], ICL vision model has recently shown potential for creating universal models that can adapt to new tasks and domains by using image-label pairs as prompts to convey task-specific information. In natural image processing, ICL-based models such as LVM [7], Painter [49], and SegGPT [50] have demonstrated strong versatility across diverse tasks.
@@ 3 Method
ICL models achieve universality by learning task-specific processing from contexts. However, standard vision ICL models require high memory due to the simultaneous processing of 3D targets and contexts. To mitigate this, we propose a strategy that adaptively balances sequential and parallel context computation, significantly reducing memory demands while maintaining consistent outputs and training gradients. Furthermore, we introduce an optimized loss function to enhance model focus on challenging segmentation regions and brain anatomy.
@@ 3.1 Model
As illustrated in Figure 2, the model contains two 3D U-Net branches: a target branch for extracting target image representations and a context branch for extracting context representations. These branches communicate through the U-shape fusion strategy at each block. The context branch receives concatenated image-label pairs as input and uses adaptive parallel-sequential processing to compute the mean context representation, which is then passed to the target branch decoder for final prediction generation.
Context Partition. Given the task-specific context $S=\{(x_{i},y_{i})\}_{i=1}^{L}$, where each $(x_{i},y_{i})$ represents an image-label pair within the context set of size $L$. After context partition, the context $S$ is divided into $n=\lceil\frac{L}{\ell}\rceil$ disjoint mini-contexts $\{s_{j}\}_{j=1}^{n}$, each containing $\ell$ image-label pairs, where $\ell$ is determined by the available memory. Larger values of $\ell$ decrease inference time but increase memory usage, and vice versa. If $L$ is not divisible by $\ell$, the final mini-context will contain fewer than $\ell$ pairs, ensuring that $\bigcup_{j=1}^{n}s_{j}=S$.
Adaptive Parallel-Sequential Processing. Adaptive Parallel-Sequential Processing (APSP) accommodates arbitrary context partitions. Specifically, each mini-context is initially processed in parallel, and the resulting context representations are subsequently fused during sequential processing.
During parallel processing, all image-label pairs within a mini-context $s_{j}$ are processed simultaneously through shared-weight 3D U-Nets in the context branch. The context representation for each $s_{j}$ is computed as $C_{j}=g(s_{j},f_{\text{enc}}(x))$, where $x$ represents the target image, $g$ denotes the context branch, and $f_{\text{enc}}$ is the encoder of the target branch. Here, $C_{j}$ is a set comprising the representations from each stage of the context decoder, defined as $C_{j}=\{\bar{c}_{\text{dec},j}^{i}\}_{i=1}^{\mathcal{D}}$, where $\mathcal{D}$ represents the number of decoder stages. The term $\bar{c}_{\text{dec},j}^{i}$ is computed via a parallel mean aggregation of image-label pairs inside $s_{j}$:
@@
The sequential computation procedure for $\bar{C}$ is outlined in Algorithm 1. The model processes each mini-context sequentially, updating $\bar{C}$ iteratively. The sequential mean operation ensures equal weighting of image-label pairs across the entire context set, making $\bar{C}$ invariant to ordering of mini-contexts. After $\bar{C}$ is obtained, and the final prediction is computed as $\hat{y}=f_{\text{dec}}(f_{\text{enc}}(x),\bar{C})$, where $f_{\text{dec}}$ denotes the decoder of the target branch. The combination of parallel and sequential mean operations in APSP guarantees consistent mean context representation computation for any context order or partition.
To save memory, intermediate features from processed mini-contexts are discarded after contributing to $\bar{C}$, with only the last mini-context retained for gradient computation during training. This reduces memory usage to the final mini-context rather than the entire context set. Moreover, when training, we employ random shuffling of the mini-contexts at each step and scale the value of $\bar{c}_{\text{dec},n}^{i}$ by a factor of $n$ during gradient computation, ensuring that the expected value of the gradients computed using APSP is equivalent to the expectation when no mini-contexts are discarded, as demonstrated by:
@@
where $\nabla\mathcal{L}_{\text{full}}$ and $\nabla\mathcal{L}_{\text{APSP-scaled},\pi}$ denote the gradients obtained from full-context processing (without mini-context exclusion) and APSP, respectively. $\pi$ represents the index of the mini-context selected for gradient computation in APSP. A proof of this equivalence is provided in the supplemental section A.1.
U-Shape Fusion Strategy. The U-shaped design makes feature propagation through two phases: first, from the target branch to the context branch at each encoder stage via Target-to-Context Fusion modules, and then from the context branch back to the target branch at each decoder stage through Context-to-Target Fusion modules, as illustrated in fig. 2. This U-shaped representation flow is actually essential, as it allows the model to avoid storing features from previously processed mini-contexts, a limitation of alternative fusion strategies. Further explanation are provided in the supplemental section A.2.
@@
where $\hat{t}^{i}$ denotes the fused target representation which is sent to the next decoder stage, and $\bar{c}^{i}$ represents the mean context representation.
@@ 3.2 Loss Function
where $\tau$ denotes the sampled tasks, and $\lambda^{\tau}\in\mathbb{R}^{+}$ represents the task-specific weighting coefficient. The procedure begins by sampling a task $\tau$ from all tasks, followed by sampling a target image $x^{\tau}$, corresponding ground truth $y^{\tau}$ and a support set $S^{\tau}$ within this task. The task-specific loss function $\mathcal{L}^{\tau}$ is then applied to compute the loss between the $\hat{y}^{\tau}$ and $y^{\tau}$.
For both segmentation and generation tasks, we employ an $L_{1}$-based loss function, with modifications to accommodate the characteristics of 3D neuroimages. Specifically, 3D neuroimage segmentation, compared to its 2D counterpart, suffers from greater sparsity of certain target structures (e.g. hippocampus, amygdala) due to the additional dimension, leading to a more severe class imbalance problem. To address this, we propose a modified $\mathcal{L}_{\text{smooth}-L_{1}}$ loss defined as:
@@
This formulation, employing a higher exponent than the standard $\mathcal{L}_{\text{smooth}-L_{1}}$, shifts the model’s focus from large, simple structures to smaller, complex regions, balancing performance across diverse segmentation tasks. We opted against Dice or Focal loss to avoid further complicating the balance between segmentation and generation tasks.
For generation tasks, we apply a standard $\mathcal{L}_{\text{smooth}-L_{1}}$ loss. Furthermore, to achieve better generative performance on brain images filled with complex anatomical structures, we incorporate an additional $\mathcal{L}_{\text{smooth}-L_{1}}$ loss on the intensity difference of the image, inspired by [38]:
@@ 4.1 Data
Datasets. To ensure robust cross-center generalization and diversity, we collected 19 datasets with multiple modalities and centers, comprising 43,674 scans. The dataset includes common neuroimaging modalities including T1, T2, FLAIR, MRA, DWI, ADC, PD, and CT. Data from 15 datasets, including 38,126 scans and 5,632 segmentation masks, were used for training and validation [54, 4, 21, 35, 3, 24, 42, 15, 1, 31, 5, 23, 17, 45, 48, 44], with a random 9:1 split. Data from the remaining 4 datasets, including 5,548 images and 1,096 segmentation masks, were used as held-out datasets [33, 39, 2, 43] to assess performance on unseen centers. The held-out dataset was divided into an 8:2 split for the meta context set (from which ICL models select context) and the test set. During training, we used both the original images and images aligned to the MNI 152 template space [16] to increase data diversity. We utilized FreeSurfer [14] to generate additional anatomical segmentation labels for 3 datasets [17, 45, 2].
@@
Neuroimaging Tasks. With these datasets, our model was trained on multiple segmentation tasks, including anatomical segmentation [8], tumor segmentation [44], vessel segmentation [54], and generation tasks, including bias field correction [18], inpainting [40], super-resolution [34], Gaussian noise removal, salt-and-pepper noise removal [34], 2D-to-3D generation [29], modality transformation [46], and skull stripping [25]. For segmentation tasks, binary masks were produced by thresholding the model outputs at 0.5, ensuring balanced results.
Image and Task Augmentation. During training, we applied data augmentation across two dimensions: image and task. In addition to standard image augmentations, we enhanced the contrast diversity of images using a randomly initialized convolutional network [47]. For task augmentation, we introduced random task overlapping and also followed the method introduced in [13]. Synthetic brain data with random modalities [9] was also added. Further details are provided in supplemental section B.2.
Detailed information on datasets and tasks is provided in the supplemental section B.

### v2

# Neuroverse3D: Developing In-Context Learning Universal Model for Neuroimaging in 3D

@@ Abstract
In-context learning (ICL), a type of universal model, demonstrates exceptional generalization across a wide range of tasks without retraining by leveraging task-specific guidance from context, making it particularly effective for the intricate demands of neuroimaging. However, current ICL models, limited to 2D inputs and thus exhibiting suboptimal performance, struggle to extend to 3D inputs due to the high memory demands of ICL. In this regard, we introduce Neuroverse3D, an ICL model capable of performing multiple neuroimaging tasks in 3D (e.g., segmentation, denoising, inpainting). Neuroverse3D overcomes the large memory consumption associated with 3D inputs through adaptive parallel-sequential context processing and a U-shaped fusion strategy, allowing it to handle an unlimited number of context images. Additionally, we propose an optimized loss function to balance multi-task training and enhance focus on anatomical boundaries. Our study incorporates 43,674 3D multi-modal scans from 19 neuroimaging datasets and evaluates Neuroverse3D on 14 diverse tasks using held-out test sets. The results demonstrate that Neuroverse3D significantly outperforms existing ICL models and closely matches task-specific models, enabling flexible adaptation to medical center variations without retraining. The code and model weights are publicly available at https://github.com/jiesihu/Neuroverse3D.
11footnotetext: Equal contribution.22footnotetext: Corresponding author: Hanyang Peng, Ting Ma.
@@ 1 Introduction
Computational neuroimage analysis has significantly advanced our understanding of the brain and non-invasive diagnostics, crucial for quantitative analysis and precision medicine. Recently, universal models trained on multi-domain datasets have gained attention for their ability to handle diverse tasks (e.g., segmentation, denoising, inpainting) and modalities (e.g., MRI-T1w, MRI-T2w, CT), demonstrating adaptability to domain shifts and even unseen tasks [12, 38, 43, 57]. In-context learning (ICL) has emerged as a promising paradigm [10, 51, 7], enabling models to adapt to domains without retraining by using image-label pairs as task-specific guidance. These informative contexts make ICL models particularly effective for segmenting tissues with intricate morphology and unifying diverse image generation tasks.
@@
While ICL models show clinical potential in neuroimaging [12, 14, 54], they are limited by dimensionality. Existing models process 3D volumes as 2D slices, which results in the loss of inter-slice correlations and volumetric context. This lack of global spatial awareness hinders performance on neuroimaging tasks, such as hippocampus segmentation which requires 3D anatomical understanding [13, 20].
@@
To address this limitations, it is crucial to develop an ICL model that captures 3D global information. However, creating a universal ICL model with 3D input presents significant challenges, as 3D images can be over 100 times larger than their 2D slices, leading to substantial memory requirements. This results in three critical issues: (1) Context size is significantly limited, hindering performance as larger contexts size demonstrably improve outputs [12, 7, 14, 52]. (2) Training models with more parameters becomes infeasible, restricting their capacity and potential performance. (3) Processing high-resolution 3D inputs becomes difficult.
@@
Figure 1: (a) Parallel processing of all contexts, which is applied by most existing ICL models. (b) Sequential processing of each context, which greatly reduces memory usage. (c) It shows the memory increment when increasing the context size. Our proposed model processes contexts in an adaptive mode, allowing for flexible parallel and sequential processing of contexts.
@@
To develop 3D universal models in neuroimaging, we introduce Neuroverse3D, an in-context learning model that takes 3D neuroimaging as input. To our knowledge, it is the first 3D ICL universal model for neuroimaging. To mitigate the significant memory demands, we propose Adaptive Parallel-Sequential Processing (APSP) and a U-shape fusion strategy to partition the entire context into multiple mini-contexts for processing, while ensuring the consistency of the model’s gradient when training and output when inferencing. As shown in Figure 1, this approach can significantly reduce memory consumption, even accommodating unlimited context sizes. Additionally, to address task imbalance from increased dimensionality and enhance anatomical boundary focus, we propose an optimized loss function, further improving performance. As the first 3D ICL model in neuroimaging, we evaluate Neuroverse3D’s potential across diverse tasks on held-out datasets. Our contributions are summarized as follows:
  • We mitigate the high memory demands of 3D image contexts using a novel APSP approach with a U-shaped fusion strategy, enabling unlimited context processing for both training and inference. This significantly improves the feasibility of deploying 3D neuroimaging ICL models in clinical practice.
@@
  • We develop an optimized loss function to address class imbalance resulting from increased dimensionality and enhance focus on challenging segmentation regions and anatomical boundaries, further improving performance.
  • We propose Neuroverse3D, the first 3D ICL universal model for neuroimaging, trained on extensive multi-center datasets. Neuroverse3D significantly outperforms other ICL models, achieving over 20 absolute Dice point gains across four tasks and matching task-specific performance, offering valuable insights for universal model development.
@@ 2.1 Domain Challenges in Medical Imaging
Deep learning models in medical imaging frequently face domain shifts due to heterogeneous imaging data distributions, which reduce model performance on new, unseen domains. Solutions like domain adaptation [59, 55, 29, 53] require target-domain fine-tuning, limiting practical use due to the need for deep learning expertise. Domain generalization aims to produce models that generalize to new domains without fine-tuning [30, 32, 58, 49]. While it alleviates the effects of domain shift, generalization remains challenging because models lack knowledge about unseen domains during training. Different from domain adaptation and domain generalization methods, universal models have emerged as a promising solution by learning generalizable image representations from a large number of datasets.
@@ 2.2 Universal Models in Medical Imaging
Universal models in medical imaging can be classified into three categories based on their prompt types. The first category utilizes symbolic prompts, such as points or bounding boxes, as demonstrated by models like SAM [43, 11, 34]. The second category employs natural language prompts [38, 39]. The third category, also known as ICL models, uses image-label pairs as prompts [14, 12, 54, 28]. These universal models, trained on extensive datasets, exhibit exceptional generalization capabilities. Among these, image-label pairs provide rich contextual information, enabling ICL models to perform accurate segmentation on unseen tasks and even handle image generation tasks. However, the richness of these prompts in ICL models’ inputs significantly increases memory consumption, a critical challenge that this study seeks to overcome.
@@ 2.3 In-Context Learning
Originally developed in natural language processing [10], ICL vision models has recently shown potential for creating universal models that can adapt to new tasks and domains by using image-label pairs as prompts to convey task-specific information. In natural image processing, ICL-based models such as LVM [7], Painter [51], and SegGPT [52] have demonstrated strong versatility across diverse tasks.
@@ 3 Method
In-Context Learning (ICL) models for neuroimaging achieve universality by learning task-specific processing from context examples, typically image-label pairs. In the following sections, we detail our Adaptive Parallel-Sequential Processing (APSP) and U-shaped fusion strategy for efficient context processing, followed by our optimized loss function for further performance improvements.
@@ 3.1 Model
As illustrated in Figure 2, the model contains two 3D U-Net branches: a target branch for extracting target image representations and a context branch for extracting context representations. These branches communicate through the U-shape fusion strategy at each stage. The context branch receives concatenated image-label pairs as input and uses APSP to compute the mean context representation, which is then passed to the target branch decoder for final prediction generation.
Context Partition. Given a task-specific context $S=\{(x_{i},y_{i})\}_{i=1}^{L}$, where each $(x_{i},y_{i})$ represents an image-label pair within a context set of size $L$, we perform context partitioning. This divides $S$ into $n=\lceil\frac{L}{\ell}\rceil$ disjoint mini-contexts $\{s_{j}\}_{j=1}^{n}$, each containing $\ell$ image-label pairs, with $\ell$ determined by available memory resources. Larger values of $\ell$ decrease inference time but increase memory usage, and vice versa. If $L$ is not divisible by $\ell$, the final mini-context will contain fewer than $\ell$ pairs, ensuring that $\bigcup_{j=1}^{n}s_{j}=S$.
Adaptive Parallel-Sequential Processing. APSP accommodates arbitrary context partitions. Specifically, each mini-context is initially processed in parallel, and the resulting context representations are subsequently fused during sequential processing.
During parallel processing, all image-label pairs within a mini-context $s_{j}$ are processed simultaneously through shared-weight 3D U-Nets in the context branch. The context representation for each mini-context $s_{j}$ is computed as $C_{j}=g(s_{j},f_{\text{enc}}(x))$, where $x$ represents the target image, $g$ denotes the context branch, and $f_{\text{enc}}$ is the encoder of the target branch. Here, $C_{j}$ is a set comprising the representations from each stage of the context decoder, defined as $C_{j}=\{\bar{c}_{\text{dec},j}^{i}\}_{i=1}^{\mathcal{D}}$, where $\mathcal{D}$ represents the number of decoder stages. $\bar{c}_{\text{dec},j}^{i}$ is computed via a parallel mean aggregation of image-label pairs inside $s_{j}$:
@@
The sequential computation procedure for $\bar{C}$ is outlined in Algorithm 1. The model processes each mini-context sequentially, updating $\bar{C}$ iteratively. The sequential mean operation ensures equal weighting of image-label pairs across the entire context set, making $\bar{C}$ invariant to the ordering of mini-contexts. After $\bar{C}$ is obtained, and the final prediction is computed as $\hat{y}=f_{\text{dec}}(f_{\text{enc}}(x),\bar{C})$, where $f_{\text{dec}}$ denotes the decoder of the target branch. The combination of parallel and sequential mean operations in APSP guarantees consistent mean context representation computation for any context order or partition.
To save memory, intermediate features from processed mini-contexts are discarded after contributing to $\bar{C}$, with only the last mini-context retained for gradient computation during training. This reduces memory usage to the final mini-context rather than the entire context set. Moreover, during training, we randomly shuffle mini-contexts and scale the last mini-context’s representation $C_{n}$ by a factor of $n$ during gradient computation. This ensures the expected value of the gradients computed using APSP is equivalent to the expectation when no mini-contexts are discarded, as demonstrated below:
@@
where $\nabla\mathcal{L}_{\text{full}}$ and $\nabla\mathcal{L}_{\text{APSP-scaled},\pi}$ denote the gradients obtained from full-context processing and APSP, respectively. $\pi$ represents the original index of the mini-context selected for gradient computation in APSP. A proof of this equivalence is provided in the supplemental Sec. A.1.
U-Shape Fusion Strategy. The U-shaped design makes feature propagation through two phases: first, from the target branch to the context branch at each encoder stage via Target-to-Context Fusion modules, and then from the context branch back to the target branch at each decoder stage through Context-to-Target Fusion modules, as illustrated in Fig. 2. This U-shaped representation flow is actually essential, as it allows the model to avoid storing features from previously processed mini-contexts, a limitation of alternative fusion strategies. Further explanation are provided in the supplemental Sec. A.2.
@@
where $\hat{t}^{i}$ denotes the fused target representation that is sent to the next decoder stage, and $\bar{c}^{i}$ represents the mean context representation.
@@ 3.2 Loss Function
where $\tau$ denotes the sampled tasks, and $\lambda^{\tau}\in\mathbb{R}^{+}$ represents the task-specific weighting coefficient. The procedure begins by sampling a task $\tau$ from all tasks, followed by sampling a target image $x^{\tau}$, corresponding ground truth $y^{\tau}$ and a support set $S^{\tau}$ within that task. The task-specific loss function $\mathcal{L}^{\tau}$ is then applied to compute the loss between the $\hat{y}^{\tau}$ and $y^{\tau}$.
For both segmentation and generation tasks, we employ an $L_{1}$-based loss function, with modifications to accommodate the characteristics of 3D neuroimages.
Specifically, 3D neuroimage segmentation, unlike its 2D counterpart, exhibits increased sparsity in structures like the hippocampus and amygdala due to the added dimension, resulting in greater class imbalance. To mitigate this, we propose a modified $\mathcal{L}_{\text{smooth}-L_{1}}$ loss:
@@
This formulation, using a higher exponent than standard $\mathcal{L}_{\text{smooth}-L_{1}}$, prioritizes challenging anatomical regions over simpler ones, balancing performance across diverse tasks. We opted against Dice or Focal loss to avoid further complicating the balance between segmentation and generation tasks.
For generation tasks, we apply a standard $\mathcal{L}_{\text{smooth}-L_{1}}$ loss. Furthermore, to achieve better generative performance on brain images filled with intricate anatomical boundaries, we incorporate an additional $\mathcal{L}_{\text{smooth}-L_{1}}$ loss on the intensity difference of the image, inspired by [40]:
@@ 4.1 Data
Datasets. To ensure robust cross-center generalization and diversity, we collected 19 datasets with multiple modalities and centers, comprising 43,674 scans. The dataset includes common neuroimaging modalities including T1, T2, FLAIR, MRA, DWI, ADC, PD, and CT. Data from 15 datasets, including 38,126 scans and 5,632 segmentation masks, were used for training and validation [56, 4, 23, 37, 3, 26, 44, 16, 1, 33, 5, 25, 18, 47, 50, 46], with a random 9:1 split. Data from the remaining 4 datasets, including 5,548 images and 1,096 segmentation masks, was used as held-out datasets [35, 41, 2, 45] to assess performance on unseen centers. The held-out dataset was divided into an 8:2 split for the meta context set (from which ICL models select context) and the test set. During training, we used both the original images and images aligned to the MNI 152 template space [17] to increase data diversity. We utilized FreeSurfer [15] to generate additional anatomical segmentation labels for 3 datasets [18, 47, 2].
@@
Neuroimaging Tasks. With these datasets, our model was trained on multiple segmentation tasks, including anatomical segmentation [8], tumor segmentation [46], vessel segmentation [56], and generation tasks, including bias field correction [19], inpainting [42], super-resolution [36], Gaussian noise removal, salt-and-pepper noise removal [36], 2D-to-3D reconstruction [31], modality transformation [48], and skull stripping [27]. For segmentation tasks, binary masks were produced by thresholding the model outputs at 0.5.
Image and Task Augmentation. During training, we applied data augmentation across two dimensions: image and task. In addition to standard image augmentations, we enhanced the contrast diversity of images using a randomly initialized convolutional network [49]. For task augmentation, we introduced random task overlapping and also followed the method introduced in [14]. Synthetic brain data with random modalities [9] was also added. Further details are provided in supplemental Sec. B.3.
Detailed information on datasets and tasks is provided in the supplemental Sec. B.